# Decoupled Generative Modeling for Human-Object Interaction Synthesis

### v1

@@ 1 Introduction
Realistic human-object interaction synthesis (HOI) is a fundamental task with broad impact in computer vision and robotics [bhatnagar2022behave, xu2025interact, liu2022hoi4d, cao2023detecting, chen2023detecting]. These capabilities form the basis of modern 3D systems, enabling human motion animation and humanoid control [zhu2023human, xue2025human, liu2025core4d]. Synthesizing interaction requires perceiving the pose of the object and awareness of the target goal, followed by the generation of a plausible sequence of human joint configurations that performs the required manipulation safely and intentionally [antoun2023human, yang2024lemon, yang2024f, xu2025intermimic]. Recent approaches [li2024controllable, wu2025human] rely on interpretable natural language instructions to specify tasks, yet producing interactions that are faithful to the prompt while keeping realism and diversity remains a difficult challenge.
@@
Prior methods [li2024controllable, wu2025human] generate 4D motion sequences of humans and objects by conditioning a diffusion model on path waypoints and text instructions. While these methods provide a strong foundation for interaction modeling that is capable of producing desired behaviors, this design introduces two challenges: i) reliance on specified intermediate waypoints that the human must follow, and ii) high optimization complexity from assigning the entire objective to a single denoising network. Specifically, for the first challenge, at inference time, the user must manually provide not only the start and goal points but also intermediate waypoints. Such reliance on externally supplied constraints narrows the model’s generative scope and reduces autonomy, while introducing procedural burden for the user. For the second challenge, HOI requires the model to simultaneously resolve the motion of both the human and the object in every frame. Assigning all objectives to a single network elevates optimization complexity and risks practical intractability. The resulting complexity often produces unsynchronized interactions, such as hovering or penetration of objects, which harms plausibility and degrades overall quality. These observations motivate us to design a new intermediate representation that is simple, flexible, informative, and still enables interactions that remain faithful to the instruction and plausible for HOI generation.
@@
Another key factor for human-level interaction is precise hand and foot control [hao2024hand, mao2022contact]. To improve the robustness of hand and foot contact and reduce undesired penetration, we introduce adversarial training applied to the object and to the distal joints. A compact discriminator focuses on hand-object interaction signals and foot dynamics, shaping the learning distribution and enabling realistic control.
For applications, the previous model [li2024controllable] enables 3D scene-aware long-sequence interaction synthesis but remains restricted to static environments. To overcome this limitation, our framework models a moving counterpart and equips the human agent to either avoid the counterpart or wait, allowing it to respond to dynamic environments. This improves practicality in dynamic scenes by aligning actions with the plan and adapting to moving obstacles.
To validate the effectiveness of DecHOI, we evaluated our method on the FullBodyManipulation [li2023object] and observed superior performance on most quantitative metrics and qualitative assessments. We also report results on unseen objects from the 3D-FUTURE [fu20213d], demonstrating robustness and generalizability. For long-sequence interaction synthesis, we present scenarios in indoor environments furnished with diverse objects, where the human agent interacts with a moving counterpart and reaches various goal points by either avoiding the counterpart or waiting. The summary of our contributions is as follows.
@@ 2.1 Human Action Generation
Human action generation aims to synthesize realistic 3D motion sequences from conditioning signals such as text descriptions, action labels, and joint-level constraints [meng2025rethinking, zhang2025energymogen, fan2024freemotion, zou2024parco, jiang2023motiongpt]. T2M [guo2022generating] samples a motion length from text and then synthesizes the motion with a temporal VAE over motion snippet codes. MDM [tevet2023human] adapts classifier-free diffusion [ho2022classifier] with a transformer backbone for high-fidelity motion, and PriorMDM [shafir2024human] treats a pre-trained diffusion model as a generative prior to enable controllable composition. ACMDM [meng2025absolute] uses absolute global joint coordinates in a streamlined transformer diffusion model, improving fidelity and text alignment. OmniControl [xie2024omnicontrol] adds spatial constraints for any joint at any time using analytic guidance with a learned realism prior, while MoMask [guo2024momask] enables text-guided temporal inpainting via masked generation with hierarchical motion tokens. However, these works primarily address single human motion and neither manipulate objects nor perceive or adapt to the surrounding scene [wang2025hsi, wang2024move, li2024genzi]. In contrast, we target scene-aware human-object interaction that avoids obstacles and manipulates objects in accordance with the specified intent.
@@ 2.2 Human-Object Interaction Synthesis
Human-object interaction synthesis (HOI) takes text instructions or object state as input and generates purposeful 3D human-object motion [xue2025guiding, jia2025primhoi, zhang2025interactanything, song2024hoianimator]. For example, OMOMO [li2023object] conditions on full per-frame object motion and enforces contact by integrating hand kinematics during whole body generation. CHOIS [li2024controllable] and HOIFHLI [wu2025human] show that per-frame object conditioning is unnecessary because language instruction paired with sparse 3D object waypoints, together with the initial human and object states, stabilizes intended behaviors. As a complementary effort, CG-HOI [diller2024cg] jointly generates human and object motion with contact cues, using contact-guided conditioning to improve realism. Despite these gains, a single network that models trajectory, pose, and contact is still hard to optimize and often produces desynchronization and contact artifacts. We instead decouple trajectory and action generation and apply contact-aware adversarial training, yielding robust synchronized HOI.
@@ 2.3 Path-conditioned HOI
Path-conditioned HOI first specifies a long-term trajectory via waypoints, object paths, or scene-level guidance, and then synthesizes it at the motion level [yi2024generating, cen2024generating, lee2023locomotion]. CHOIS generates synchronized long-horizon human-object motion in 3D scenes by conditioning diffusion on sparse object waypoints obtained from Habitat [savva2019habitat, szot2021habitat] for the given start and target positions. NIFTY [kulkarni2024nifty] couples an object-conditioned motion diffusion model with a learned interaction field that evaluates distances to a feasible interaction manifold and guides sampling toward contact-consistent behaviors. OMOMO [li2023object] takes per-frame object motion data as input and synthesizes manipulation by denoising, incorporating hand motion to maintain contact during whole body generation. However, in inference, these formulations retain a fixed plan or guidance and therefore do not adapt to external scene changes. By contrast, our method updates the plan, enabling adaptation to external scene changes while preserving high-level intent.
@@ 3.1 Background: Denoising Diffusion Model
Denoising Diffusion Probabilistic Models (DDPMs) [ho2020denoising, tevet2023human] are generative models that learn to approximate complex data distributions by gradually adding noise and then removing it. This paradigm has become dominant in recent work on human action generation [tevet2023human, shafir2024human, meng2025absolute, xie2024omnicontrol] and interaction synthesis [li2023object, li2024controllable, wu2025human, xu2023interdiff]. The forward process progressively corrupts a clean sample $\mathbf{x}_{0}$ by adding Gaussian noise over $N$ steps, defined as:
@@
where $\mathbf{c}$ denotes conditioning inputs such as text embeddings, object geometry, or start and goal states. Following prior formulations [li2024controllable, wu2025human], our denoising network predicts the clean sequence rather than the noise:
@@ 3.2 Decoupled Generative Modeling
Jointly estimating and optimizing human and object poses is non-trivial for a single network to handle. We address this challenge with decoupled generative modeling that separates trajectory generation from action generation. At the start of the pipeline the inputs are the object pose sequence $P_{o}\in\mathbb{R}^{T\times 12}$ with $T$ frames, where each frame contains the global 3D position and a 9-parameter relative rotation matrix, and the human pose sequence $P_{h}\in\mathbb{R}^{T\times D_{h}}$, where $D_{h}$ comprises global joint coordinates and per-joint 6D rotation parameters. To ensure accurate interaction and instruction following, inputs also include object geometry $B\in\mathbb{R}^{1024\times 3}$ represented as a Basis Point Set (BPS) [prokudin2019efficient] and a text instruction.
@@
Trajectory Generator. We formulate the trajectory generator (TG) as a conditional denoising diffusion model [ho2020denoising]. Given an input sequence, we apply a forward noising process for $N$ steps following a Markov chain. For conditioning, the human and object poses of the start frame remain clean, and the object position in the end frame is also kept clean to set the goal point. Then these noisy data are concatenated with the geometry embedding $F_{\text{obj}}\in\mathbb{R}^{D}$, computed by a simple MLP encoder from $B$. We obtain the conditioning input $X\in\mathbb{R}^{T\times(12+D_{h}+D)}$ for the denoising network. A transformer-based denoising network of TG generates both global object and human trajectories that are faithful to the instruction. For example, given the instruction “Lift the chair, move the chair, and put down the chair,” the object trajectory rises in the early part of the sequence and descends as it approaches the goal point. To learn this alignment, we incorporate a text embedding $F_{\text{text}}\in\mathbb{R}^{D}$ obtained from a CLIP encoder [radford2021learning] into the model inputs. In contrast to prior methods [li2024controllable, wu2025human] that condition by concatenating the embedding along the sequence dimension, we use a cross-attention layer to apply the conditioning explicitly, which yields stronger alignment between text and motion features. Afterward, the TG yields continuous 3D paths for the object $\hat{\mathcal{T}}_{o}\in\mathbb{R}^{T\times 3}$ and the human $\hat{\mathcal{T}}_{h}\in\mathbb{R}^{T\times 3}$.
@@
Action Generator. Once reliable human and object trajectories have been generated, we use them as richer conditions. The action generator (AG) receives the same input $X$ as the trajectory generator (TG), except that the noisy global positions of the object and the human root in all $T$ frames are replaced with the trajectories produced by TG. This dense conditioning provides stronger priors and reduces the complexity of the learning objective. AG is a lightweight diffusion model that employs a transformer-based denoising network with the same architecture as TG. Given the conditioned noisy input, it generates the full pose of the object $\hat{P}_{o}\in\mathbb{R}^{T\times 12}$ and the pose of the human joints $\hat{P}_{h}\in\mathbb{R}^{T\times D_{h}}$. The human pose generated is then used for parametric human modeling, and we apply SMPL-X [pavlakos2019expressive] to reconstruct the human mesh and pose.
@@ 3.3 Adversarial Training for Distal Joints
Realistic human-object interaction (HOI) requires not only plausible motion but also reliable contact [cseke2025pico, gu2024contactgen, jiang2024scaling]. Language instructions such as “Lift the chair” or “Kick the box” rely on distal joints in the hands and feet. Existing models often cause the hands and feet to drift into empty space or intersect objects [yuan2023physdiff, xu2025interact, mao2022contact]. To address these issues, we introduce an adversarial training mechanism that regularizes contact by focusing on the distal joints.
In general, the most reliable cue to decide whether an input is real or fake is the distance between the object surface and the distal joints [mao2022contact, cseke2025pico, huang2022capturing, tripathi2023deco]. In ground truth, these distances are small due to complete contact, whereas generated results often exhibit larger gaps. Optimizing the generator so that the discriminator cannot distinguish real from fake therefore acts as a regularizer that drives contact toward completeness. Motivated by this observation, we design a compact discriminator $\mathcal{D}$ that differentiates real and fake using distal joint kinematics together with object geometry. As shown in Fig. 3, $\mathcal{D}$ receives global coordinates of the hands $H\in\mathbb{R}^{T\times 2\times 3}$ and feet $F\in\mathbb{R}^{T\times 2\times 3}$ over all $T$ frames, obtained by forward kinematics from the 6D joint rotations. In addition, we sample $M$ points from the geometry of the object $B$ to obtain $B’\in\mathbb{R}^{T\times M\times 3}$ and apply the relative rotation and translation for each frame. For real data, we feed clean inputs $x$, and for fake data, we use the outputs $\hat{x}$ produced by AG. The discriminator is trained with the following loss:
@@
Lin-OMOMO [li2023object] - - 10.04 0.44 0.45 16.81 6.65 0.75 0.56 0.59 0.70 0.65 17.45 29.20 75.65 -
Pred-OMOMO [li2023object] 2.34 9.66 6.99 0.38 0.63 13.01 6.68 0.67 0.48 0.52 0.59 0.60 23.63 49.12 51.86 1.23
CHOIS [li2024controllable] 1.92 8.01 6.15 0.41 0.68 1.58 8.31 0.74 0.53 0.58 0.66 0.61 18.86 44.04 51.86 1.23
HOIFHLI [wu2025human] 1.73 7.65 4.77 0.38 0.62 2.06 8.55 0.77 0.61 0.64 0.58 0.61 19.31 40.87 50.96 1.18
@@
Table 1: Quantitative comparison on the FullBodyManipulation [li2023object] with CHOIS [li2024controllable], HOIFHLI [wu2025human], and OMOMO [li2023object] variants (Lin-OMOMO and Pred-OMOMO) across four categories of evaluation metrics. Arrows indicate direction: ($\uparrow$) means higher is better, ($\downarrow$) means lower is better, and ($\rightarrow$) means closer to the real data value is better. The real-data DIVDIV reference is 9.02.
@@
Lin-OMOMO [li2023object] - - 7.92 0.48 0.57 8.85 0.26 0.20
Pred-OMOMO [li2023object] 4.72 10.92 6.61 0.45 0.59 5.16 0.40 0.17
CHOIS [li2024controllable] 5.75 10.28 4.20 0.42 0.61 2.04 0.46 0.18
@@
Table 2: Quantitative results on the 3D-FUTURE [fu20213d]. DecHOI achieves better trajectory accuracy, motion stability, and contact realism than CHOIS [li2024controllable] and OMOMO [li2023object] baselines.

### v2

@@ 1 Introduction
Realistic human-object interaction synthesis (HOI) is a fundamental task with broad impact in computer vision and robotics [2, 45, 27, 3, 5]. These capabilities form the basis of modern 3D systems, enabling human motion animation and humanoid control [56, 48, 26]. Synthesizing interaction requires perceiving the pose of the object and awareness of the target goal, followed by the generation of a plausible sequence of human joint configurations that performs the required manipulation safely and intentionally [1, 51, 50, 47]. Recent approaches [23, 43] rely on interpretable natural language instructions to specify tasks, yet producing interactions that are faithful to the prompt while keeping realism and diversity remains a difficult challenge.
@@
Prior methods [23, 43] generate 4D motion sequences of humans and objects by conditioning a diffusion model on path waypoints and text instructions. While these methods provide a strong foundation for interaction modeling that is capable of producing desired behaviors, this design introduces two challenges: i) reliance on specified intermediate waypoints that the human must follow, and ii) high optimization complexity from assigning the entire objective to a single denoising network. Specifically, for the first challenge, at inference time, the user must manually provide not only the start and goal points but also intermediate waypoints. Such reliance on externally supplied constraints narrows the model’s generative scope and reduces autonomy, while introducing procedural burden for the user. For the second challenge, HOI requires the model to simultaneously resolve the motion of both the human and the object in every frame. Assigning all objectives to a single network elevates optimization complexity and risks practical intractability. The resulting complexity often produces unsynchronized interactions, such as hovering or penetration of objects, which harms plausibility and degrades overall quality. These observations motivate us to design a new intermediate representation that is simple, flexible, informative, and still enables interactions that remain faithful to the instruction and plausible for HOI generation.
@@
Another key factor for human-level interaction is precise hand and foot control [13, 28]. To improve the robustness of hand and foot contact and reduce undesired penetration, we introduce adversarial training applied to the object and to the distal joints. A compact discriminator focuses on hand-object interaction signals and foot dynamics, shaping the learning distribution and enabling realistic control.
For applications, the previous model [23] enables 3D scene-aware long-sequence interaction synthesis but remains restricted to static environments. To overcome this limitation, our framework models a moving counterpart and equips the human agent to either avoid the counterpart or wait, allowing it to respond to dynamic environments. This improves practicality in dynamic scenes by aligning actions with the plan and adapting to moving obstacles.
To validate the effectiveness of DecHOI, we evaluated our method on the FullBodyManipulation [24] and observed superior performance on most quantitative metrics and qualitative assessments. We also report results on unseen objects from the 3D-FUTURE [9], demonstrating robustness and generalizability. For long-sequence interaction synthesis, we present scenarios in indoor environments furnished with diverse objects, where the human agent interacts with a moving counterpart and reaches various goal points by either avoiding the counterpart or waiting. The summary of our contributions is as follows.
@@ 2.1 Human Action Generation
Human action generation aims to synthesize realistic 3D motion sequences from conditioning signals such as text descriptions, action labels, and joint-level constraints [30, 54, 8, 57, 19]. T2M [12] samples a motion length from text and then synthesizes the motion with a temporal VAE over motion snippet codes. MDM [39] adapts classifier-free diffusion [16] with a transformer backbone for high-fidelity motion, and PriorMDM [36] treats a pre-trained diffusion model as a generative prior to enable controllable composition. ACMDM [29] uses absolute global joint coordinates in a streamlined transformer diffusion model, improving fidelity and text alignment. OmniControl [44] adds spatial constraints for any joint at any time using analytic guidance with a learned realism prior, while MoMask [11] enables text-guided temporal inpainting via masked generation with hierarchical motion tokens. However, these works primarily address single human motion and neither manipulate objects nor perceive or adapt to the surrounding scene [41, 42, 25]. In contrast, we target scene-aware human-object interaction that avoids obstacles and manipulates objects in accordance with the specified intent.
@@ 2.2 Human-Object Interaction Synthesis
Human-object interaction synthesis (HOI) takes text instructions or object state as input and generates purposeful 3D human-object motion [49, 18, 55, 37]. For example, OMOMO [24] conditions on full per-frame object motion and enforces contact by integrating hand kinematics during whole body generation. CHOIS [23] and HOIFHLI [43] show that per-frame object conditioning is unnecessary because language instruction paired with sparse 3D object waypoints, together with the initial human and object states, stabilizes intended behaviors. As a complementary effort, CG-HOI [7] jointly generates human and object motion with contact cues, using contact-guided conditioning to improve realism. Despite these gains, a single network that models trajectory, pose, and contact is still hard to optimize and often produces desynchronization and contact artifacts. We instead decouple trajectory and action generation and apply contact-aware adversarial training, yielding robust synchronized HOI.
@@ 2.3 Path-conditioned HOI
Path-conditioned HOI first specifies a long-term trajectory via waypoints, object paths, or scene-level guidance, and then synthesizes it at the motion level [52, 4, 22]. CHOIS generates synchronized long-horizon human-object motion in 3D scenes by conditioning diffusion on sparse object waypoints obtained from Habitat [35, 38] for the given start and target positions. NIFTY [21] couples an object-conditioned motion diffusion model with a learned interaction field that evaluates distances to a feasible interaction manifold and guides sampling toward contact-consistent behaviors. OMOMO [24] takes per-frame object motion data as input and synthesizes manipulation by denoising, incorporating hand motion to maintain contact during whole body generation. However, in inference, these formulations retain a fixed plan or guidance and therefore do not adapt to external scene changes. By contrast, our method updates the plan, enabling adaptation to external scene changes while preserving high-level intent.
@@ 3.1 Background: Denoising Diffusion Model
Denoising Diffusion Probabilistic Models (DDPMs) [15, 39] are generative models that learn to approximate complex data distributions by gradually adding noise and then removing it. This paradigm has become dominant in recent work on human action generation [39, 36, 29, 44] and interaction synthesis [24, 23, 43, 46]. The forward process progressively corrupts a clean sample $\mathbf{x}_{0}$ by adding Gaussian noise over $N$ steps, defined as:
@@
where $\mathbf{c}$ denotes conditioning inputs such as text embeddings, object geometry, or start and goal states. Following prior formulations [23, 43], our denoising network predicts the clean sequence rather than the noise:
@@ 3.2 Decoupled Generative Modeling
Jointly estimating and optimizing human and object poses is non-trivial for a single network to handle. We address this challenge with decoupled generative modeling that separates trajectory generation from action generation. At the start of the pipeline the inputs are the object pose sequence $P_{o}\in\mathbb{R}^{T\times 12}$ with $T$ frames, where each frame contains the global 3D position and a 9-parameter relative rotation matrix, and the human pose sequence $P_{h}\in\mathbb{R}^{T\times D_{h}}$, where $D_{h}$ comprises global joint coordinates and per-joint 6D rotation parameters. To ensure accurate interaction and instruction following, inputs also include object geometry $B\in\mathbb{R}^{1024\times 3}$ represented as a Basis Point Set (BPS) [33] and a text instruction.
@@
Trajectory Generator. We formulate the trajectory generator (TG) as a conditional denoising diffusion model [15]. Given an input sequence, we apply a forward noising process for $N$ steps following a Markov chain. For conditioning, the human and object poses of the start frame remain clean, and the object position in the end frame is also kept clean to set the goal point. Then these noisy data are concatenated with the geometry embedding $F_{\text{obj}}\in\mathbb{R}^{D}$, computed by a simple MLP encoder from $B$. We obtain the conditioning input $X\in\mathbb{R}^{T\times(12+D_{h}+D)}$ for the denoising network. A transformer-based denoising network of TG generates both global object and human trajectories that are faithful to the instruction. For example, given the instruction “Lift the chair, move the chair, and put down the chair,” the object trajectory rises in the early part of the sequence and descends as it approaches the goal point. To learn this alignment, we incorporate a text embedding $F_{\text{text}}\in\mathbb{R}^{D}$ obtained from a CLIP encoder [34] into the model inputs. In contrast to prior methods [23, 43] that condition by concatenating the embedding along the sequence dimension, we use a cross-attention layer to apply the conditioning explicitly, which yields stronger alignment between text and motion features. Afterward, the TG yields continuous 3D paths for the object $\hat{\mathcal{T}}_{o}\in\mathbb{R}^{T\times 3}$ and the human $\hat{\mathcal{T}}_{h}\in\mathbb{R}^{T\times 3}$.
@@
Action Generator. Once reliable human and object trajectories have been generated, we use them as richer conditions. The action generator (AG) receives the same input $X$ as the trajectory generator (TG), except that the noisy global positions of the object and the human root in all $T$ frames are replaced with the trajectories produced by TG. This dense conditioning provides stronger priors and reduces the complexity of the learning objective. AG is a lightweight diffusion model that employs a transformer-based denoising network with the same architecture as TG. Given the conditioned noisy input, it generates the full pose of the object $\hat{P}_{o}\in\mathbb{R}^{T\times 12}$ and the pose of the human joints $\hat{P}_{h}\in\mathbb{R}^{T\times D_{h}}$. The human pose generated is then used for parametric human modeling, and we apply SMPL-X [32] to reconstruct the human mesh and pose.
@@ 3.3 Adversarial Training for Distal Joints
Realistic human-object interaction (HOI) requires not only plausible motion but also reliable contact [6, 10, 20]. Language instructions such as “Lift the chair” or “Kick the box” rely on distal joints in the hands and feet. Existing models often cause the hands and feet to drift into empty space or intersect objects [53, 45, 28]. To address these issues, we introduce an adversarial training mechanism that regularizes contact by focusing on the distal joints.
In general, the most reliable cue to decide whether an input is real or fake is the distance between the object surface and the distal joints [28, 6, 17, 40]. In ground truth, these distances are small due to complete contact, whereas generated results often exhibit larger gaps. Optimizing the generator so that the discriminator cannot distinguish real from fake therefore acts as a regularizer that drives contact toward completeness. Motivated by this observation, we design a compact discriminator $\mathcal{D}$ that differentiates real and fake using distal joint kinematics together with object geometry. As shown in Fig. 3, $\mathcal{D}$ receives global coordinates of the hands $H\in\mathbb{R}^{T\times 2\times 3}$ and feet $F\in\mathbb{R}^{T\times 2\times 3}$ over all $T$ frames, obtained by forward kinematics from the 6D joint rotations. In addition, we sample $M$ points from the geometry of the object $B$ to obtain $B’\in\mathbb{R}^{T\times M\times 3}$ and apply the relative rotation and translation for each frame. For real data, we feed clean inputs $x$, and for fake data, we use the outputs $\hat{x}$ produced by AG. The discriminator is trained with the following loss:
@@
Lin-OMOMO [24] - - 10.04 0.44 0.45 16.81 6.65 0.75 0.56 0.59 0.70 0.65 17.45 29.20 75.65 -
Pred-OMOMO [24] 2.34 9.66 6.99 0.38 0.63 13.01 6.68 0.67 0.48 0.52 0.59 0.60 23.63 49.12 51.86 1.23
CHOIS [23] 1.92 8.01 6.15 0.41 0.68 1.58 8.31 0.74 0.53 0.58 0.66 0.61 18.86 44.04 51.86 1.23
HOIFHLI [43] 1.73 7.65 4.77 0.38 0.62 2.06 8.55 0.77 0.61 0.64 0.58 0.61 19.31 40.87 50.96 1.18
@@
Table 1: Quantitative comparison on the FullBodyManipulation [24] with CHOIS [23], HOIFHLI [43], and OMOMO [24] variants (Lin-OMOMO and Pred-OMOMO) across four categories of evaluation metrics. Arrows indicate direction: ($\uparrow$) means higher is better, ($\downarrow$) means lower is better, and ($\rightarrow$) means closer to the real data value is better. The real-data DIVDIV reference is 9.02.
@@
Lin-OMOMO [24] - - 7.92 0.48 0.57 8.85 0.26 0.20
Pred-OMOMO [24] 4.72 10.92 6.61 0.45 0.59 5.16 0.40 0.17
CHOIS [23] 5.75 10.28 4.20 0.42 0.61 2.04 0.46 0.18
@@
Table 2: Quantitative results on the 3D-FUTURE [9]. DecHOI achieves better trajectory accuracy, motion stability, and contact realism than CHOIS [23] and OMOMO [24] baselines.